the length of a rectangle is 7 meters less than twice he width. If the perimeter is 40 meters, find the dimensions of the rectangle. Answer by mananth(16949) 
 Perimeter = 40 feet
Length 1 ( 2 x -7 )
Width = x ft
Perimeter =2(L+W)
Perimeter = 2(((2x-7)+x)
Perimeter = 2((2x-7 +x)
Perimeter =4x-14+2x
40 =6x-14
40 -=6x
40+14=6 x
54=6x
x=9 ft Width
Length =11ft